# Drift2Matrix: Kernel-Induced Self Representation for Concept Drift Adaptation in Co-evolving Time Series

### v2

@@ 4.2 Adaptation to Concept Drift
For $b$ sliding windows $\{W_{1},\ldots,W_{b}\}$ [^2] The window size is determined through a heuristic method that balances the granularity of concept identification with the representational capacity of Drift2Matrix. For details, see Sec. 6.1 and Appendix D., Drift2Matrix constructs individual kernel representations for each window. Let’s consider $k$ distinct concepts identified across these windows, denoted as $\mathbf{C}_{c\in[1,k]}=\{\mathbf{C}_{1},\cdots,\mathbf{C}_{k}\}$. It is important to know that the concepts identified from the subseries $\mathbf{S}_{p}$ within the $p$-$th$ $(p\in[1,b])$ window may differ from those in other windows. This reveals the variety of concepts in time series and the demand for a dynamic representation.
@@
$P(\mathbf{C}r\rightarrow\mathbf{C}m|W_{p}\rightarrow W_{p+1},S_{i})=\frac{\sum_{\zeta}\Psi_{p,p+1}^{r,\zeta}\Lambda_{p,p+1}^{\zeta,m}}{\sum_{\zeta_{1}}\sum_{\zeta_{2}}\Psi_{p,p+1}^{\zeta_{1},\zeta_{2}}\Lambda_{p,p+1}^{\zeta_{1},\zeta_{2}}}$ (4)
@@
$\Psi_{p,p+1}^{r,m}=$ $\frac{\eta\left(\mathbf{C}_{r}\rightarrow\mathbf{C}_{m}|\mathcal{T}r\left(S_{i}|W_{p}\right)\right)}{|\mathcal{T}r(S_{i}|W_{p})|},$ (5)
$\Lambda_{p,p+1}^{r,m}=$ $\sum_{l=1}^{p-1}\frac{\min\{\eta(\mathbf{C}_{r},W_{l}),\eta(\mathbf{C}_{m},W_{l+1})\}}{\max\{\eta(\mathbf{C}_{r},W_{l}),\eta(\mathbf{C}_{m},W_{l+1})\}}$
@@ 4.3 Integration into Deep Learning Backbones
One of the key strengths of Drift2Matrix is its flexibility, which allows it to be easily integrated into most modern deep learning backbones. Here, we take Autoencoder-Drift2Matrix (Auto-D2M) as an example, which comprises an Encoder, a Kernel Representation Layer, and a Decoder.
@@
Decoder: The decoder reconstructs the input from the refined latent representations $\hat{\mathbf{S}}_{\Theta_{d}}=\text{Decoder}_{\Theta_{d}}(\hat{\mathbf{H}}_{\Theta_{e}})$, where $\hat{\mathbf{S}}_{\Theta_{d}}$ represents the reconstructed time series segments.
@@
$\mathcal{L}(\Theta)=\frac{1}{2}\|\mathbf{S}-\hat{\mathbf{S}}_{\Theta_{d}}\|_{F}^{2}+\lambda_{1}\|\mathbf{\Theta_{s}}\|_{1}+\lambda_{2}\|\Phi(\mathbf{H}_{\Theta_{e}})-\Phi(\mathbf{H}_{\Theta_{e}})\mathbf{\Theta_{s}}\|_{F}^{2},$
where $\mathbf{\Theta}=\{\mathbf{\Theta_{e}},\mathbf{\Theta_{s}},\mathbf{\Theta_{d}}\}$ includes all learnable parameters, with $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ balancing the different loss components. Specifically, $\lambda_{1}$ promotes sparsity in the self-representation $\mathbf{\Theta_{s}}$, and $\lambda_{2}$ preserves the self-representation property.
@@ 5.1 Behavior of the Representation Matrix
The core of Drift2Matrix is the kernel representation matrix $\mathbf{Z}$, which encapsulates the relationships and concepts within time series. Without loss of generality, let $\mathbf{S}=[\mathbf{S}^{(1)},\mathbf{S}^{(2)},\cdots,\mathbf{S}^{(k)}]$ be ordered according to their concept. Ideally, we wish to obtain a representation $\mathbf{Z}$ such that each point is represented as a combination of points belonging to the same concept, i.e., $\mathbf{S}^{(i)}=\mathbf{S}^{(i)}\mathbf{Z}^{(i)}$. In this case, $\mathbf{Z}$ in Eq. 1 has the $k$-block diagonal structure (up to permutations), i.e.,
@@
Theorem 5.1
In Drift2Matrix, the representation matrix obtained for a permuted input data is equivalent to the permutation-transformed original representation matrix. Let $\mathbf{Z}$ be feasible to $\Phi(\mathbf{S})=\Phi(\mathbf{S})\mathbf{Z}$, then $\tilde{\mathbf{Z}}=\mathbf{P^{T}ZP}$ is feasible to $\Phi(\tilde{\mathbf{S}})=\Phi(\tilde{\mathbf{S}})\tilde{\mathbf{Z}}$. (See Appendix C.1 for proof.)
@@ 5.2 Kernel-Induced Representation
The kernel-induced representation in Drift2Matrix is a key step that reveals nonlinear relationships among co-evolving time series in a high-dimensional space. This approach transforms complex, intertwined patterns in the original time series, which may not be discernible in low-dimensional spaces, into linearly separable entities in the transformed space. Essentially, it allows for a deeper and more nuanced understanding of the dynamics hidden within complex time series structures. Through kernel transformation, previously obscured correlations and patterns become discernible, enabling more precise and insightful analysis of co-evolving time series data.
@@
Theorem 5.2
Drift2Matrix reveals nonlinear relationships among series in a high-dimensional space while simultaneously preserving the local manifold structure of series. (See Appendix C.2 for proof.)
@@ 6 Experiments
This section presents our experiments to evaluate the effectiveness of Drift2Matrix. The experiments were designed to answer the following questions:
(Q1) Effectiveness: How well does Drift2Matrix identify and track concept drift?
(Q2) Accuracy: How accurately does Drift2Matrix forecast future concept and series value?
(Q3) Scalability: How does Drift2Matrix perform in online forecasting scenarios?
@@ 6.1 Data and experimental setup
The data utilized in our experiments consists of a synthetic dataset (SyD) constructed to allow the controllability of the structures/numbers of concepts and the availability of ground truth, as well as several real-life datasets: GoogleTrend Music Player dataset (MSP), Customer electricity load (ELD) data, Chlorine concentration data (CCD), Earthquake data (EQD), Electrooculography signal (EOG), Rock dataset (RDS), two financial datasets (Stock1 & Stock2), four ETT (ETTh1, ETTh2, ETTm1, ETTm2), Traffic and Weather datasets. In our kernel representation learning process, we chose the Gaussian kernel function. Detailed information about these datasets and the setup of the kernel function can be found in Appendix F. The source code is public to the research community333 https://anonymous.4open.science/r/Drift2Matrix-main-86B7.
In the learning process, a fixed window slides over all the series and generates subseries under different windows. Then, we learn a kernel representation for the subseries in each window. Given our focus on identifying concepts and concept drift, varying window sizes actually demonstrate Drift2Matrix’s ability to extend across multi-scale time series. Specifically, smaller window sizes represent a low-scale perspective, uncovering short-term subtle concept variations (fluctuations), while larger window sizes (high-scale) reflect overall concept trends (moderation). In this paper, we employ MDL (Minimum Description Length) techniques Rissanen (1998) on the segment-score obtained through our kernel-induced representation to determine a window size  [^4] For ETT, Traffic, and Weather datasets, we consider sizes of 96, 192, 336, and 720, following the standard settings used by most methods. that establishes highly similar or repetitive concepts across different windows (see Appendix D). Although adaptively determining a domain-agnostic window size forms a part of our work, to ensure fairness in comparison, all models, including Drift2Matrix, were evaluated using the same window size settings in our experiments.
@@ 6.2 Q1: Effectiveness
Drift2Matrix’s forecasting effectiveness is evaluated through its ability to identify important concepts. Due to space limitations, here we only describe our results for the SyD and Stock1 datasets, the outputs with the other datasets are shown in Appendix H.2. Our method for automatically estimating the optimal size of the sliding windows to obtain suitable concepts is detailed in Appendix D. The method allows obtaining window sizes of 78 and 17 for Synthetic and Stock1, respectively (see Fig 5). The Drift2Matrix output of SyD has already been presented in Fig. 1 of Sec. 1. As already seen, our method automatically captures typical concepts in a given co-evolving time series (Fig. 1 (b-c)), and dynamic concept drifts (Fig. 1 (d)). Drift2Matrix views co-evolving time series as an ecosystem, enabling precise detection of interconnected dynamics and drifts. This also facilitates forecasting future concepts and values (The grey areas in Fig. 1 (d)).
@@
Fig. 2(a) illustrates various concepts identified in the Stock1 dataset, revealing patterns of market volatility. These discovered concepts are meaningful as they encapsulate different market behaviors, such as steady trends, spikes, or dips, corresponding to distinct phases in market activity. For example, Concept 4 represents stable periods with low volatility, while Concepts 2 and 5 capture moments of sudden market fluctuations or high-risk events. Additionally, the subtle differences between Concepts 2 and 5 highlight the model’s ability to detect gradual drifts in market behavior. Using these concepts, we trace series exhibiting concept drift over different windows. For the purpose of illustration, in Fig. 2(b), we plot the heatmap of each concept across different windows and their corresponding 2D visualizations using t-Distributed Stochastic Neighbor Embedding (t-SNE) Van der Maaten & Hinton (2008). Darker heatmap cells indicate more prevalent concepts. The exclusive appearance of $\mathbf{C}_{2}$ and $\mathbf{C}_{5}$ in the 6th window, absent in the preceding ones, denotes significant changes and fluctuations in the financial markets – i.e., possibly induced by the COVID-19 pandemic. Fig. 2(c) shows the concept drifts and predictions for three random stocks (NASDAQ: MRK, MU and ETSY) from Stock1, demonstrating Drift2Matrix’s ability to track and forecast concept drifts.
6.3 Q2: Accuracy
For real datasets, we lack the ground truth for validating the obtained concepts. Instead, we validate the value and gain of the discovered concepts for time series forecasting as they are employed in the forecasting formula Eq. 6. In this section, we evaluate the forecasting performance of the proposed model against seventeen different models, utilizing the Root Mean Square Error (RMSE) as an evaluative metric. Due to space limitations, we only present results for seven comparison models here; the complete experimental results can be found in the Appendix H.3. These seven models include four forecasting models (ARIMA Box (2013), KNNR Chen & Paschalidis (2019), INFORMER Zhou et al. (2021), and a ensemble model N-BEATS Oreshkin et al. (2019)), and three are concept-drift models (Cogra Miyaguchi & Kajino (2019), OneNet Wen et al. (2024) and OrBitMap Matsubara & Sakurai (2019)). For the existing methods, we use the codes released by the authors, and the details of the parameter settings can be found in Appendix G.
Table 1: Models’ forecasting performance, in terms of RMSE
@@ 6.3 Q2: Accuracy
ARIMA KNNR INFORMER N-BEATS Cogra OneNet OrbitMap Drift2Matrix Auto-D2M
@@
Stock1 (×$(\times 10^{-2})$ 17 $2.635$ $2.348$ $2.127$ $1.035$ $2.137$ $0.923$ $1.003$ $\mathbf{0.878}$ 0.902
Stock2 (×$(\times 10^{-2})$ 11 $2.918$ $2.761$ $1.064$ $0.607$ $1.367$ $0.312$ $0.747$ $\mathbf{0.303}$ 0.317
@@
720 1.096 0.917 0.938 0.867 0.863 0.833 0.841 0.840 0.832
While forecasting series task is not our main focus, we provide a comparison of Drift2Matrix with other models. Results for the extended Auto-D2M, are included but not part of the comparison. Complete experimental results can be found in the Appendix H.3.
Table 1 shows the forecasting performance of the models. We see that our model consistently outperforms the other models, achieving the lowest forecasting error on most datasets. ARIMA has the ability to capture seasonality patterns within time series; however, when the various seasonalities are noncontiguous, the models face difficulties in capturing complex, nonlinear dynamic interactions between time series. N-BEATS, a state-of-the-art deep learning model, while generally effective due to its ensemble-based architecture, does not consistently perform as well as Drift2Matrix or OneNet, particularly in capturing concept drift across multiple time series. OneNet, like N-BEATS, achieves good results due to its ensemble-based strengths. However, Drift2Matrix achieves comparable results. Notably, Drift2Matrix is not primarily designed as a forecasting model; rather, it focuses on uncovering concepts and tracking concept drift. This unique focus allows Drift2Matrix to excel in identifying complex, dynamic interactions within time series data. Meanwhile, OrbitMap, while also concept-aware, is hindered by its necessity for predefined concepts and struggles with handling multiple time series.
@@ 6.4 Q3: Scalability
To further illustrate the predictive scalability of Drift2Matrix, we employed it for one of the most challenging tasks in time series analysis – i.e., online forecasting, leveraging the discovered concepts. For this task, our objective is to forecast upcoming unknown future events, at any given moment, while discarding redundant information. This approach is inherently aligned with online learning paradigms, where the model continually learns and adapts to new data points, making it highly pertinent in the dynamic landscape of financial markets. We conducted tests on the Stock2 dataset.
Fig. 3 illustrates the online forecasting examples on four stocks (NASDAQ: CSX, ULTA, UNP and BK) and showcases snapshots at several time-stamps. The original data at the top of Fig. 3(a) elucidates the daily volatility fluctuations for these four stocks. The lower part of Fig. 3(a) unveils the outcomes for online forecasting, showcasing how our model anticipates series behavior over time. For stock ULTA, depicted by the green line, all the compared models, including our model, encounter challenges in predicting the abnormal behavior of the first high volatility (Fig. 3(b-1)) due to the absence of antecedent knowledge. However, post encountering this anomalous behavior, our model accurately anticipates the timing of the second anomalous high-volatility behavior (Fig. 3(b-2)) and the ensuing volatility behavior, attributing to Drift2Matrix’s ability to model concept drift by leveraging the interrelations among multiple time series. Specifically, for a single-series concept drift model, it is impossible to predict the second anomalous behavior accurately if there is no periodic pattern in it; whereas for our model, since our concept drift is based on correlations between series, when other series start to show some anomalous volatility behavior (albeit small), our model is also able to predict the next volatility of multiple time series in a holistic way.

### v3

@@ 4.2 Adaptation to Concept Drift
For $b$ sliding windows $\{W_{1},\ldots,W_{b}\}$ [^2] Based on the learned $\mathbf{Z}$, the window size is determined through a heuristic method that balances the granularity of concept identification with the representational capacity of CORAL. For details, see Sec. 6.1 and Appendix D., CORAL constructs individual kernel representations for each window. Let’s consider $k$ distinct concepts identified across these windows, denoted as $\mathbf{C}_{c\in[1,k]}=\{\mathbf{C}_{1},\cdots,\mathbf{C}_{k}\}$. It is important to know that the concepts identified from the subseries $\mathbf{S}_{p}$ within the $p$-$th$ $(p\in[1,b])$ window may differ from those in other windows. This reveals the variety of concepts in time series and the demand for a dynamic representation.
@@
[…]
@@
[…]
[yshift=1em]rightnode1Transition probability of $\mathbf{C}_{r}$ to $\mathbf{C}_{m}$ in $S_{i}$annotate[yshift=-1em]below,rightnode2Transition probability of $\mathbf{C}_{r}$ to $\mathbf{C}_{m}$ in Ecosystem $\mathbf{S}$
@@ 4.3 Integration into Deep Learning Backbones
One of the key strengths of CORAL is its flexibility, which allows it to be easily integrated into most modern deep learning backbones. Here, we take Autoencoder-CORAL (Auto-CORAL) as an example, which comprises an Encoder, a Kernel Representation Layer, and a Decoder.
@@
Decoder: The decoder reconstructs the input from the refined latent representations $\hat{\mathbf{S}}_{\Theta_{d}}=\text{Decoder}_{\Theta_{d}}(\hat{\mathbf{H}}_{\Theta_{e}})$, where $\hat{\mathbf{S}}_{\Theta_{d}}$ represents the reconstructed series segments.
@@
$\mathcal{L}_{(\Theta)}$ $=\frac{1}{2}\|\mathbf{S}-\hat{\mathbf{S}}_{\Theta_{d}}\|_{F}^{2}+\lambda_{1}\|\mathbf{\Theta_{s}}\|_{1}$
$\quad+\lambda_{2}\|\Phi(\mathbf{H}_{\Theta_{e}})-\Phi(\mathbf{H}_{\Theta_{e}})\mathbf{\Theta_{s}}\|_{F}^{2}$
where $\mathbf{\Theta}=\{\mathbf{\Theta_{e}},\mathbf{\Theta_{s}},\mathbf{\Theta_{d}}\}$ includes learnable parameters, with $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ balancing the different loss components. Specifically, $\lambda_{1}$ promotes sparsity in the self-representation $\mathbf{\Theta_{s}}$, and $\lambda_{2}$ preserves the self-representation property.
@@ 5.1 Behavior of the Representation Matrix
The core of CORAL is the kernel representation matrix $\mathbf{Z}$, which encapsulates the relationships and concepts within time series. Without loss of generality, let $\mathbf{S}=[\mathbf{S}^{(1)},\mathbf{S}^{(2)},\cdots,\mathbf{S}^{(k)}]$ be ordered according to their concept. Ideally, we wish to obtain a representation $\mathbf{Z}$ such that each point is represented as a combination of points belonging to the same concept, i.e., $\mathbf{S}^{(i)}=\mathbf{S}^{(i)}\mathbf{Z}^{(i)}$. In this case, $\mathbf{Z}$ in Eq. 1 has the $k$-block diagonal structure (up to permutations), i.e.,
@@
Theorem 5.1.
In CORAL, the representation matrix obtained for a permuted input data is equivalent to the permutation-transformed original representation matrix. Let $\mathbf{Z}$ be feasible to $\Phi(\mathbf{S})=\Phi(\mathbf{S})\mathbf{Z}$, then $\tilde{\mathbf{Z}}=\mathbf{P^{T}ZP}$ is feasible to $\Phi(\tilde{\mathbf{S}})=\Phi(\tilde{\mathbf{S}})\tilde{\mathbf{Z}}$. (See Appendix C.2 for proof.)
@@ 5.2 Kernel-Induced Representation
The kernel-induced representation in CORAL is a key step that reveals nonlinear relationships among co-evolving time series in a high-dimensional space. This approach transforms complex, intertwined patterns in the original time series, which may not be discernible in low-dimensional spaces, into linearly separable entities in the transformed space. Essentially, it allows for a deeper and more nuanced understanding of the dynamics hidden within complex time series structures. Through kernel transformation, previously obscured correlations and patterns become discernible, enabling more precise and insightful analysis of co-evolving time series data.
@@
Theorem 5.2.
CORAL reveals nonlinear relationships among series in a high-dimensional space while simultaneously preserving the local manifold structure of series. (See Appendix C.3 for proof.)
@@ 6 Experiments
This section presents our experiments to evaluate the effectiveness of CORAL. The experiments were designed to answer the following questions:
(Q1) Effectiveness: How well does CORAL identify and track concept drift?
(Q2) Accuracy: How accurately does CORAL forecast future concept?
(Q3) Scalability: How does CORAL perform in online scenarios?
@@ 6.1 Data and experimental setup
The data utilized in our experiments consists of a synthetic dataset (SyD) constructed to allow the controllability of the structures/numbers of concepts and the availability of ground truth, as well as several real-life datasets: GoogleTrend Music Player dataset (MSP), Customer electricity load (ELD) data, Chlorine concentration data (CCD), Earthquake data (EQD), Electrooculography signal (EOG), Rock dataset (RDS), two financial datasets (Stock1 & Stock2), four ETT (ETTh1, ETTh2, ETTm1, ETTm2), Traffic and Weather datasets. In our kernel representation learning process, we chose the Gaussian kernel function. Detailed information about these datasets and the setup of the kernel function can be found in Appendix F.
In the learning process, a fixed window slides over all the series and generates subseries under different windows. Then, we learn a kernel representation for the subseries in each window. Given our focus on identifying concepts and concept drift, varying window sizes actually demonstrate CORAL’s ability to extend across multi-scale time series. Specifically, smaller window sizes represent a low-scale perspective, uncovering short-term subtle concept variations (fluctuations), while larger window sizes (high-scale) reflect overall concept trends (moderation). In this paper, we employ MDL (Minimum Description Length) techniques (Rissanen, 1998) on the segment-score obtained through our kernel-induced representation to determine a window size  [^3] For ETT, Traffic, and Weather datasets, we consider sizes of 96, 192, 336, and 720, following the standard settings used by most methods. that establishes highly similar or repetitive concepts across different windows (see Appendix D). Although adaptively determining a domain-agnostic window size forms a part of our work, to ensure fairness in comparison, all models, including CORAL, were evaluated using the same window size settings in our experiments.
@@ 6.2 Q1: Effectiveness
CORAL’s forecasting effectiveness is evaluated through its ability to identify important concepts. Due to space limitations, here we only describe our results for the SyD and Stock1 datasets, the outputs with the other datasets are shown in Appendix G.2. Our method for automatically estimating the optimal size of the sliding windows to obtain suitable concepts is detailed in Appendix D. The method allows obtaining window sizes of 78 and 17 for Synthetic and Stock1, respectively (see Fig 6). The CORAL output of SyD has already been presented in Fig. 1 of Sec. 1. As already seen, our method automatically captures typical concepts in a given co-evolving time series (Fig. 1 (b-c)), and dynamic concept drifts (Fig. 1 (d)). CORAL views co-evolving time series as an ecosystem, enabling precise detection of interconnected dynamics and drifts. This also facilitates forecasting future concepts and values (The grey areas in Fig. 1 (d)).
@@
Fig. 2(a) illustrates various concepts identified in the Stock1 dataset, revealing patterns of market volatility. These discovered concepts are meaningful as they encapsulate different market behaviors, such as steady trends, spikes, or dips, corresponding to distinct phases in market activity. For example, Concept 4 represents stable periods with low volatility, while Concepts 2 and 5 capture moments of sudden market fluctuations or high-risk events. Additionally, the subtle differences between Concepts 2 and 5 highlight the model’s ability to detect gradual drifts in market behavior. Using these concepts, we trace series exhibiting concept drift over different windows. For the purpose of illustration, in Fig. 2(b), we plot the heatmap of each concept across different windows and their corresponding 2D visualizations using t-Distributed Stochastic Neighbor Embedding (t-SNE) (Van der Maaten & Hinton, 2008). Darker heatmap cells indicate more prevalent concepts. The exclusive appearance of $\mathbf{C}_{2}$ and $\mathbf{C}_{5}$ in the 6th window, absent in the preceding ones, denotes significant changes and fluctuations in the financial markets – i.e., possibly induced by the COVID-19 pandemic. Fig. 2(c) shows the concept drifts and predictions for three random stocks (NASDAQ: MRK, MU and ETSY) from Stock1, demonstrating CORAL’s ability to track and forecast concept drifts.
Table 1: Models’ forecasting performance, in terms of RMSE. The best results are in red color and the second best in blue color.
@@
ARIMA KNNR INFORMER N-BEATS Cogra OneNet OrbitMap CORAL Auto-CORAL
@@
Stock1 (×$(\times 10^{-2})$ 17 2.635 2.348 2.127 1.035 2.137 0.923 1.003 0.878 0.902
Stock2 (×$(\times 10^{-2})$ 11 2.918 2.761 1.064 0.607 1.367 0.312 0.747 0.303 0.317
@@
720 1.096 0.917 0.938 0.867 0.863 0.833 0.841 0.840 0.839
While forecasting task is not our main focus, we provide a comparison of CORAL with other models. Results for the extended Auto-CORAL. Complete experimental results can be found in the Appendix G.3, Table 7.
6.3 Q2: Accuracy
For real datasets, we lack the ground truth for validating the obtained concepts. Instead, we validate the value and gain of the discovered concepts for time series forecasting as they are employed in Eq. 6. Here, we evaluate the forecasting performance of the proposed model against seventeen different models, utilizing the Root Mean Square Error (RMSE) as an evaluative metric. Due to space limitations, we only present results for seven comparison models here; the complete experimental results can be found in the Appendix G.3, Table 7. These seven models include four forecasting models (ARIMA (Box, 2013), KNNR (Chen & Paschalidis, 2019), INFORMER (Zhou et al., 2021), and a ensemble model N-BEATS (Oreshkin et al., 2019)), and three are concept-drift models (Cogra (Miyaguchi & Kajino, 2019), OneNet (Wen et al., 2024) and OrBitMap (Matsubara & Sakurai, 2019)). For existing methods, we use the codes released by the authors, and the details of the parameter settings can be found in Appendix F.
Table 1 presents the forecasting performance of various models. Our model consistently outperforms most baselines, achieving the lowest forecasting error across multiple datasets. While state-of-the-art deep learning models such as N-BEATS and OneNet perform well due to their ensemble-based architecture, CORAL achieves comparable results as a single model. Notably, CORAL is not primarily designed for forecasting; instead, it focuses on uncovering concepts and tracking concept drift. This unique capability enables it to effectively capture complex, dynamic interactions within time series data. In contrast, OrbitMap, despite being concept-aware, relies on predefined concepts and struggles with multiple time series, limiting its adaptability. Additionally, Fig. 3 provides visualizations of N-BEATS and CORAL on the ETTh1 dataset, demonstrating CORAL’s strengths in handling co-evolving time series with concept drift.
Figure 3: Visualization of N-BEATS and CORAL on ETTh1.
@@ 6.4 Q3: Scalability
To further illustrate the scalability of CORAL, we employed it for one of the most challenging tasks in time series analysis – i.e., online forecasting, leveraging the discovered concepts. For this task, our objective is to forecast upcoming unknown future events, at any given moment, while discarding redundant information. This approach is inherently aligned with online learning paradigms, where the model continually learns and adapts to new data points, making it highly pertinent in the dynamic landscape of financial markets. We conducted tests on the Stock2 dataset.
Fig. 4 illustrates the online forecasting examples on four stocks (NASDAQ: CSX, ULTA, UNP and BK) at different time-stamps. The original data at the top of Fig. 4(a) depicts daily volatility fluctuations, while the lower part illustrates our model’s online forecasting performance. For stock ULTA (green line), all models, including ours, struggle to predict the first high-volatility event (Fig.4(b-1)) due to the lack of prior information. However, after encountering this anomaly, our model successfully anticipates the second high-volatility event (Fig.4(b-2)) and subsequent fluctuations. This is attributed to CORAL’s ability to model concept drift by leveraging inter-series correlations. Unlike single-series models, which cannot predict such anomalies without periodic patterns, CORAL detects early volatility signals in related series, enabling holistic forecasting of multi-series volatility.